OpenStudy (anonymous):

please help with formula for q6

OpenStudy (anonymous):

1)Vo* (R3-(B*R1*R3)+R1+R2+A*R2)/(R3-B*R1*R3)=VIN 2)i1=V0/(R3-B*R1*R3) 3)vo = vth 4)Rth=Vth/Isc

OpenStudy (anonymous):

used the above, didn't work. got one chance left

OpenStudy (anonymous):

The circuit elements have the following values: Vin=15V, R1=0.5Ω, R2=1.5Ω, R3=5Ω, A=2 and B=1.5.

OpenStudy (anonymous):

I have the same values and got them all wrong using those formulas as well.

OpenStudy (anonymous):

@malchiva

OpenStudy (anonymous):

Vo = Vth = 3 I = 2.4 Rth = 4